### v2

@@ 7 Conclusion
This paper addresses the opportunities and challenges for leveraging an online advertisement system with LLM. We first discuss essential requirements which a reasonable LLM advertisement system should satisfy. To fulfill such desiderata, we introduce several frameworks and substantiate their validity. We compare proposed frameworks with their advantages and limitations. Then we discuss further technical and practical challenges in designing efficient LLM-based online advertisement system. Finally, as a usecase of independent interest, we present the potential of advanced dynamic creative optimization by using LLM.
@@ 8 Acknowledgements
This work is partially supported by DARPA QuICC, NSF AF:Small #2218678, and NSF AF:Small #2114269, and a grant from an NSF CAREER AWARD 1942230, ONR YIP award N00014-22-1-2271, ARO’s Early Career Program Award 310902-00001, Meta grant 23010098, HR00112090132 (DARPA/RED), HR001119S0026 (DARPA/GARD), Army Grant No. W911NF2120076, NIST 60NANB20D134, the NSF award CCF2212458, an Amazon Research Award and an award from Capital One. The authors thank Sébastien Lahaie for valuable discussions.
@@ B.2 Further challenges and practical concerns
Recall that our overall architecture consists of (i) modification, (ii) bidding, (iii) prediction, and (iv) auction modules. Although each of the process should be serialized in order to handle a single advertiser’s process, it is important to observe that the process (i)-(iii) is totally independent between the advertisers. Hence, one efficient to way to minimize the latency is to process steps (i) to (iii) in an asynchronous manner between the advertisers. To fit the desired latency requirement, one may consider excluding some advertisers such that the steps (i) to (iii) are not completed within some maximal delay. After collecting all the feasible parameters, LAS can pass the required parameters to step (iv) and run the auction.
@@
First, consider a simple case in which all advertisers use LAS modification and static bidding model. In this case, one can consider every combination of advertisements as a subset of $[n]$, and then generate modified output with corresponding bid. Further, if there is an advertiser with dynamic bidding, LAS can simply delegate the bid generating process to that advertiser for every combination that includes this advertiser’s ad.

### v3

@@ 7 Conclusion
This paper addresses the opportunities and challenges for leveraging an online advertisement system with LLM. We first discuss essential requirements which a reasonable LLM advertisement system should satisfy. To fulfill such desiderata, we introduce several frameworks and substantiate their validity. We compare proposed frameworks with their advantages and limitations. Then we discuss further technical and practical challenges in designing efficient online advertising with LLM. Finally, as a usecase of independent interest, we present the potential of advanced dynamic creative optimization by using LLM.
@@ 8 Acknowledgements
The work is partially supported by DARPA QuICC, ONR MURI’2024 award on Algorithms, Learning, and Game Theory, Army-Research Laboratory (ARL) #W911NF2410052, NSF AF:Small #2218678, and NSF AF:Small #2114269, and a grant from an NSF CAREER AWARD 1942230, ONR YIP award N00014-22-1-2271, ARO’s Early Career Program Award 310902-00001, Meta grant 23010098, HR00112090132 (DARPA/RED), HR001119S0026 (DARPA/GARD), Army Grant No. W911NF2120076, NIST 60NANB20D134, the NSF award CCF2212458, an Amazon Research Award and an award from Capital One. The authors thank Sébastien Lahaie for valuable discussions.
@@ B.2 Further challenges and practical concerns
Recall that our overall architecture consists of (i) modification, (ii) bidding, (iii) prediction, and (iv) auction modules. Although each of the process should be serialized in order to handle a single advertiser’s process, it is important to observe that the process (i)-(iii) is totally independent between the advertisers. Hence, one efficient to way to minimize the latency is to process steps (i) to (iii) in an asynchronous manner between the advertisers. To fit the desired latency requirement, one may consider excluding some advertisers such that the steps (i) to (iii) are not completed within some maximal delay. After collecting all the feasible parameters, LLMA can pass the required parameters to step (iv) and run the auction.
@@
First, consider a simple case in which all advertisers use LLMA modification and static bidding model. In this case, one can consider every combination of advertisements as a subset of $[n]$, and then generate modified output with corresponding bid. Further, if there is an advertiser with dynamic bidding, LLMA can simply delegate the bid generating process to that advertiser for every combination that includes this advertiser’s ad.